Find square root of 1+i


1
Expert's answer
2022-03-23T15:25:43-0400

"\sqrt{1+i}=a+bi"

"1+i=(a+bi)^2"

"1+i=(a^2-b^2)+2abi"

Equating real and imaginary parts gives us;

"1=a^2-b^2"

1=2ab

This is a simultaneous equation and so we are going to solve by substituting;

"a=\frac{1}{2b}"

"1=(\frac{1}{2b})^2-b^2=\frac{1}{4b^2}-b^2"

"4b^2=1-4b^4"

"4b^4+4b^2-1=0"

"b^2=\frac{-4 \plusmn \sqrt{4^2-4(4)(-1)}}{2(4)}=\frac{-1 \plusmn \sqrt{2}}{2}"

"b^2=\frac{-1-\sqrt{2}}{2}" has no real roots so we take the positive value

"b=\plusmn \frac{\sqrt{2(\sqrt{2}-1)}}{2}"

"a= \plusmn \frac{2\sqrt{\sqrt{2}-1}+\sqrt{2\sqrt{2}-1}}{2}"

So finally the two possible roots of the complex number 1+i are ;

"\plusmn (\frac{2\sqrt{\sqrt{2}-1}+\sqrt{2\sqrt{2}-1}}{2}+ \frac{\sqrt{2(\sqrt{2}-1)}}{2}i)"